OpenStudy (anonymous):

Which of the following is a solution of 3x2 = 7x - 3? negative 7 plus or minus the square root of 13 divided by 6 7 plus or minus the square root of 85 divided by 6 negative 7 plus or minus the square root of 85 divided by 6 x - 7 plus or minus the square root of 13 divided by 6

OpenStudy (anonymous):

i think its either A or D

OpenStudy (anonymous):

3x^2 = 7x - 3 3x^2 - 7x + 3 = 0 x^2 - 7/3x + 1 = 0 x^2 - 7/3x = - 1 x^2 - 7/3x + 49/36 = - 1 + 49/36 (x - 7/6)^2 = 13/36 x - 7/6 = ± √(13/36) x = 7/6 ± √(13)/6 x = 1/6[7 + √(13)], 1/6[7 - √(13)] answer// http://uk.answers.yahoo.com/question/index?qid=20131219071412AAe5Jbb